# IMPLICATIVE BE ALGEBRAS AS ORTHOGONALITY SPACES

### v3

@@ Abstract.
We obtain an orthogonality space by endowing an implicative involutive BE algebra with an appropriate orthogonality relation; for these spaces, we also study the particular case of implicative-orthomodular lattices. Moreover, we define the Sasaki projections as well as the commutativity relation on these structures, proving that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. We also give characterizations of implicative-Boolean algebras, and we show that the center of an implicative-orthomodular lattice is an implicative-Boolean algebra. We prove that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if it admits a full Sasaki set of projections. Finally, based on Sasaki maps on implicative involutive BE algebras, we introduce the notion of Sasaki spaces, proving that, in the case when a complete implicative involutive BE algebra admits a full Sasaki set of projections, it is a Sasaki space. We also characterize the Dacey spaces defined by implicative involutiv BE algebras.
@@ 1. Introduction
The orthomodular lattices were redefined and studied by A. Iorgulescu ([12]) based on involutive m-BE algebras, and she gave two generalizations of orthomodular lattices: orthomodular softlattices and orthomodular widelattices. Based on implicative involutive BE algebras, we redefined in [4] the orthomodular lattices, by introducing and studying the implicative-orthomodular lattices (IOML for short), and we gave certain characterizations of these structures. The distributivity of implicative-orthomodular lattices was investigated in [5]. The notion of Sasaki projections on orthocomplemented lattices was introduced by U. Sasaki ([20]), and this notion was extended to other structures ([8, 9, 3]).
In this paper, we obtain an orthogonality space by endowing an implicative involutive BE algebra with an appropriate orthogonality relation, and we investigate the properties of these spaces. Special results are proved for the particular case of implicative-orthomodular lattices as orthogonality spaces. We define and study the Sasaki projections, and based on these notions we define and characterize the commutativity relation on implicative-orthomodular lattices. Moreover, we prove that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. Based on the divisibility condition, we introduce and characterize the implicative-Boolean algebras, proving that the center of an implicative-orthomodular lattice is an implicative-Boolean algebra. We also introduce the notion of Sasaki set of projections, proving that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if it admits a full Sasaki set of projections. Inspired from [16], we define the Sasaki maps on the associated orthogonality space of an implicative involutive BE algebras and we show that, in the case of implicative-orthomodular lattices, the restriction and corestriction of a Sasaki projection is a Sasaki map. We also characterize the Dacey spaces defined by implicative involutiv BE algebras. Finally, we prove that, in the case when a complete implicative involutive BE algebra admits a full Sasaki set of projections, then it is a Sasaki space.
@@ 2. Implicative involutive BE algebras as orthogonality spaces
The implicative involutive BE algebras were introduced by A. Iorgulescu in [13], and their properties were also studied in [4].
@@ A BE algebra is called implicative if it satisfies the following condition, for all $x,y\in X$:
$(1)$-$(4)$ See [4].
@@
An implicative-orthomodular lattice (IOML for short) is an implicative involutive BE algebra satisfying any one of the equivalent conditions from Lemma 2.9.
@@
and the relation $x\vee y:=(x^{\prime}\wedge y^{\prime})^{\prime}=x^{*}\rightarrow y$. The unary operation “” is defined as $x^{*}:=x\rightarrow 0=(x\wedge 0^{\prime})^{\prime}=x^{\prime}$.
@@
3. Sasaki projections on implicative-orthomodular lattices
We introduce the notions of Sasaki projections and commutativity relations on implicative-orthomodular lattices, and we study their properties. We show that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. Moreover, the composition of two Sasaki projections $\varphi_{a}$, $\varphi_{b}$ on an implicative-orthomodular lattice is commutative if and only if $a$ commutes with $b$.
@@ 3. Sasaki projections on implicative-orthomodular lattices
$(2)$ $a\leq_{L}x$ implies $\varphi_{a}x=a;$
$(3)$ $\varphi_{\varphi_{a}x}x=\varphi_{a}x;$
$(4)$ $a\leq_{Q}x$ implies $\varphi_{x}a=a;$
$(5)$ $\varphi_{a}$ is monotone.
@@
$(1)$ if $\varphi_{a}=\varphi_{b}=\varphi_{1}$, then $\varphi_{a\Cap b}=\varphi_{1}$. $(2)$ $\varphi_{a}\varphi_{a}x=\varphi_{a}x;$
@@ 4. Implicative-Boolean center
In this section, based on the divisibility relation on implicative involutive BE algebras, we define the involutive-Boolean algebras and we give certain characterizations of these algebras. We also introduce the notion of the center $\mathcal{CC}(X)$ of an implicative-orthomodular algebra $X$, proving that $\mathcal{CC}(X)$ is an involutive-Boolean subalgebra of $X$. Moreover, we prove that an implicative involutive BE algebra $X$ is an implicative orthomodular lattice if and only if every two orthogonal elements of $X$ are contained in an implicative-Boolean subalgebra of $X$.
@@
5. Sasaki set of projections
@@ 6. Connection with Dacey and Sasaki spaces
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be its associated orthogonality space, where $X^{\prime}=X\setminus\{0\}$. If $A,B\subseteq X^{\prime}$, denote: $A^{\perp}=\{x\in X\mid x\perp a$, for all $a\in A\}$, $A\wedge B=A\cap B$, $A\vee B=(A^{\perp}\wedge B^{\perp})^{\perp}$. Clearly, $A\subseteq B$ implies $B^{\perp}\subseteq A^{\perp}$.
@@
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be its associated orthogonality space. Then $(\mathcal{C}(X^{\prime},\perp),\rightarrow,^{*},X^{\prime})$ is an implicative involutive BE algebra.
@@
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be its associated orthogonality space. The following are equivalent:
@@
If $(X,\rightarrow,^{*},1)$ be a complete implicative-orthomodular lattice, then $X$ and $\mathcal{C}(X^{\prime},\perp)$ are isomorphic as implicative-orthomodular lattice.
@@
Let $(X,\rightarrow,^{*},1)$ be a complete implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be its associated orthogonality space ($X^{\prime}=X\setminus\{0\}$). Since $X$ admits a full Sasaki set of projections, then by Theorem 5.9 it is a complete implicative-orthomodular lattice. Applying Proposition 6.11, for any orthoclosed subset $A$ of $X^{\prime}$ there exists a Sasaki map to $A$. Hence $(X^{\prime},\perp)$ is a Sasaki space. ∎
@@
Declarations
Ethical Approval
This declaration is not applicable.
Competing interests
The author has no relevant financial or non-financial interests to disclose.
Funding
No funds, grants, or other support was received.
Availability of data and materials
This declaration is not applicable.
Authors’ contributions
This declaration is not applicable.

### v4

@@ Abstract.
We obtain an orthogonality space by endowing an implicative involutive BE algebra with an appropriate orthogonality relation; for these spaces, we also study the particular case of implicative-orthomodular lattices. Moreover, we define the Sasaki projections as well as the commutativity relation on these structures, proving that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. We also give characterizations of implicative-Boolean algebras, and we show that the center of an implicative-orthomodular lattice is an implicative-Boolean algebra. We prove that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if it admits a full Sasaki set of projections. Finally, based on Sasaki maps on implicative involutive BE algebras, we introduce the notion of Sasaki spaces, proving that, in the case when a complete implicative involutive BE algebra admits a full Sasaki set of projections, it is a Sasaki space. We also characterize the Dacey spaces defined by implicative involutive BE algebras.
@@ 1. Introduction
The orthomodular lattices were redefined and studied by A. Iorgulescu ([12]) based on involutive m-BE algebras, and she gave two generalizations of orthomodular lattices: orthomodular softlattices and orthomodular widelattices. Based on implicative involutive BE algebras, we redefined in [5] the orthomodular lattices, by introducing and studying the implicative-orthomodular lattices (IOML for short), and we gave certain characterizations of these structures. The distributivity of implicative-orthomodular lattices was investigated in [6]. The notion of Sasaki projections on orthocomplemented lattices was introduced by U. Sasaki ([20]), and this notion was extended to other structures ([9, 3]). The composition of Sasaki projections on orthomodular lattices was investigated in [4].
In this paper, we obtain an orthogonality space by endowing an implicative involutive BE algebra with an appropriate orthogonality relation, and we investigate the properties of these spaces. Special results are proved for the particular case of implicative-orthomodular lattices as orthogonality spaces. We define and study the Sasaki projections, and based on these notions we define and characterize the commutativity relation on implicative-orthomodular lattices. Moreover, we prove that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. Based on the divisibility condition, we introduce and characterize the implicative-Boolean algebras, proving that the center of an implicative-orthomodular lattice is an implicative-Boolean algebra. We also introduce the notion of Sasaki set of projections, proving that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if it admits a full Sasaki set of projections. Inspired from [16], we define the Sasaki maps on the associated orthogonality space of an implicative involutive BE algebras and we show that, in the case of implicative-orthomodular lattices, the restriction and corestriction of a Sasaki projection is a Sasaki map. We also characterize the Dacey spaces defined by implicative involutive BE algebras. Finally, we prove that, in the case when a complete implicative involutive BE algebra admits a full Sasaki set of projections, then it is a Sasaki space.
@@ 2. Implicative involutive BE algebras as orthogonality spaces
The implicative involutive BE algebras were introduced by A. Iorgulescu in [13], and their properties were also studied in [24] and [5].
@@ A BE algebra is called implicative if it satisfies the following condition, for all $x,y\in X$:
$(1)$ Since $x\leq_{L}y$, we have $x^{*}=x\rightarrow y^{*}$, so that, using $(Impl)$ we get $y^{*}\rightarrow x=x^{*}\rightarrow y=(x\rightarrow y^{*})\rightarrow y=(y\rightarrow x^{*})\rightarrow y=y$, that is $y^{*}\leq_{L}x^{*}$. The converse follows similarly.
$(2)$ From $x\leq_{Q}y$ we have $x=x\Cap y$, and by $(iG)$ we get $x\rightarrow y^{*}=(x\Cap y)\rightarrow y^{*}=y\rightarrow(x\Cap y)^{*}=y\rightarrow((x^{*}\rightarrow y^{*})\rightarrow y^{*})=(x^{*}\rightarrow y^{*})\rightarrow(y\rightarrow y^{*})=(x^{*}\rightarrow y^{*})\rightarrow y^{*}=(x\Cap y)^{*}=x^{*}$. It follows that $x\leq_{L}y$.
$(3)$ Applying $(Impl)$, we have $x\rightarrow(y\rightarrow x)^{*}=(y\rightarrow x)\rightarrow x^{*}=(x^{*}\rightarrow y^{*})\rightarrow x^{*}=x^{*}$ and $x\rightarrow(x^{*}\rightarrow y)^{*}=(x^{*}\rightarrow y)\rightarrow x^{*}=x^{*}$, so that $x\leq_{L}y\rightarrow x$ and $x\leq_{L}x^{*}\rightarrow y$.
$(4)$ It follows from $x^{*}\leq_{L}x\rightarrow y^{*}$ and $y^{*}\leq_{L}x\rightarrow y^{*}$, applying $(1)$.
@@
An implicative-orthomodular lattice (IOML for short) is an implicative involutive BE algebra satisfying the equivalent conditions from Lemma 2.9.
@@
and the relation $x\vee y:=(x^{\prime}\wedge y^{\prime})^{\prime}=x^{*}\rightarrow y$. The unary operation “” is defined as $x^{*}:=x\rightarrow 0=(x\wedge 0^{\prime})^{\prime}=x^{\prime}$. Note that $\leq_{L}$ is the partial order of the lattice $(X,\wedge,\vee,^{*},0,1)$.
@@
3. Sasaki projections on implicative involutive BE algebras
We introduce the notions of Sasaki projections and commutativity relations on implicative involutive BE algebras, and we study their properties. We show that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. Moreover, the composition of two Sasaki projections $\varphi_{a}$, $\varphi_{b}$ on an implicative-orthomodular lattice is commutative if and only if $a$ commutes with $b$.
@@ 3. Sasaki projections on implicative involutive BE algebras
$(2)$ $a\leq_{L}x$ implies $\varphi_{a}x=a;$ $(3)$ $\varphi_{\varphi_{a}x}x=\varphi_{a}x;$
$(4)$ $a\leq_{Q}x$ implies $\varphi_{x}a=a;$ $(5)$ $\varphi_{a}$ is monotone.
@@
$(1)$ if $\varphi_{a}=\varphi_{b}=\varphi_{1}$, then $\varphi_{a\Cap b}=\varphi_{1};$ $(2)$ $\varphi_{a}\varphi_{a}x=\varphi_{a}x;$
@@ 4. Implicative-Boolean center
In this section, based on the divisibility relation on implicative involutive BE algebras, we define the involutive-Boolean algebras and we give certain characterizations of these algebras. We also introduce the notion of the center $\mathcal{CC}(X)$ of an implicative-orthomodular lattice $X$, proving that $\mathcal{CC}(X)$ is an involutive-Boolean subalgebra of $X$. Moreover, we prove that an implicative involutive BE algebra $X$ is an implicative orthomodular lattice if and only if every two orthogonal elements of $X$ are contained in an implicative-Boolean subalgebra of $X$.
@@
5. Sasaki set of projections on implicative involutive BE algebras
@@ 6. Connection with Dacey and Sasaki spaces
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be the associated orthogonality space, where $X^{\prime}=X\setminus\{0\}$. If $A,B\subseteq X^{\prime}$, denote: $A^{\perp}=\{x\in X\mid x\perp a$, for all $a\in A\}$, $A\wedge B=A\cap B$, $A\vee B=(A^{\perp}\wedge B^{\perp})^{\perp}$. Clearly, $A\subseteq B$ implies $B^{\perp}\subseteq A^{\perp}$.
@@
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be the associated orthogonality space. Then $(\mathcal{C}(X^{\prime},\perp),\rightarrow,^{*},X^{\prime})$ is an implicative involutive BE algebra.
@@
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be the associated orthogonality space. The following are equivalent:
@@
If $(X,\rightarrow,^{*},1)$ be a complete implicative-orthomodular lattice, then $X$ and $\mathcal{C}(X^{\prime},\perp)$ are isomorphic as implicative-orthomodular lattices.
@@
Let $(X,\rightarrow,^{*},1)$ be a complete implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be the associated orthogonality space ($X^{\prime}=X\setminus\{0\}$). Since $X$ admits a full Sasaki set of projections, then by Theorem 5.9 it is a complete implicative-orthomodular lattice. Applying Proposition 6.11, for any orthoclosed subset $A$ of $X^{\prime}$ there exists a Sasaki map to $A$. Hence $(X^{\prime},\perp)$ is a Sasaki space. ∎